[Date Prev][Date Next][Thread Prev][Thread Next][Search] [Main Index] [Thread Index] [HEASARC Mailing List Archives]

CGRO Bi-Weekly Status Report

Compton Observatory Science Report #150 Thursday, February 3, 1994
Chris Shrader, Compton Observatory Science Support Center

Questions or comments can be sent to the Compton SSC.
Phone:  301/286-8434
e-mail:   NSI_DECnet:  GROSSC::SHRADER 
          Internet: shrader@grossc.gsfc.nasa.gov

                         Spacecraft Status

The observatory and all subsystems to perform nominally.
Operational testing of the GRO Remote Tracking Station (GRTS)
continued as the TDRS-1 satellite continued its eastward drift.
TDRS-1 is expected to be on station by 02/08/94. During the last
pointing period, a 15% enhancement in the telemetry coverage was
achieved through GRTS.    

                    Science Support Center News

The CGRO fellowship applications for 1994/95 underwent peer-review
this week. It is likely that CGRO Fellowships will be awarded to
the four top ranked individuals. There was a strong consensus among
the participants in the review that the overall quality of the
applicants was quite excellent, and thus the competition intense.
Results will be announced shortly.

The CGRO Cycle-4 NRA can be retrieved electronically in Postscript
format on GROSSC (by anonymous ftp, open grossc.gsfc.nasa.gov, the
files are in /pub/nra/cycle4, by DECnet copy,
grossc::anon_dir:[nra.cycle4]*.*). Appendices A-D can also be read
on Xmosaic, and the PostScript distributions are also available on
the Goddard LHEA public ftp area on "legacy.gsfc.nasa.gov".
Questions about the NRA and/or the Cycle-4 program can be directed
to the SSC. Additional information is available on GRONEWS. 

Software for remote submission of proposal forms will be available
shortly. It will be highly familiar to ROSAT and/or ASCA proposers.
See GRONEWS for details. 

                      Guest Investigator News

In a joint PI-GI investigation, Dr. Mark Leising and collaborators
at Clemson and George Mason Universities, UNH and NRL have
requested that the observatory be repointed in response to the
transition of the X-ray binary Cygnus X-1 into a low, hard X-ray
state known as the "gamma-1" state. Theoretical calculations by Dr.
Leising and collaborators predict a possible redistribution of
photons from the x-ray domain to the MeV gamma-ray domain as a
result of suppression of Compton cooling of the inner disk.
Measurements with OSSE during the current low state (the source
typically remains in such  state for a time scale of days-weeks)
provide an excellent opportunity to test this model and associated

Dr. James Ling of the visited the GRO-SSC to discuss the eventual
export and installation of the BATSE earth-occultation package to
the SSC for general guest investigator use. This is an extensive
analysis package which has ben developed by Dr. Ling and
collaborators at JPL to utilize improved models of the earth
occultation (ingress-egress) profiles to enhance the sensitivity of
flux and spectral measurements attainable through occultation
analysis. Continuous BATSE data will be available at the SSC in the
very near future (see the BATSE report below). 

                        Instrument Reports


OSSE operations are normal.

In viewing period 316, the Z-axis target was Cen A (PI team), and
the X-axis target was CTA 102 (Key project). The Sun was accessible
along the OSSE scan plane, and slewing to the Sun in response to
BATSE triggers was enabled.

In the target-of-opportunity viewing period 318.1, the Z-axis
target is Cyg X-1 in its low state (PI team), and the X-axis
targets are NGC 7123 (PI team) and PSR 0540-693 (PI team). Solar
slewing is again enabled.

Since the last report, OSSE has slewed to the Sun on two occasions:
TJD 9378.9424 and 9384.3435.

Data from viewing periods 32 and 33 were delivered to the Compton
GRO Science Support Center Archive last week. The targets during vp
32 were NGC 3783 and the Crab Nebula and pulsar. During vp 33, the
targets were the Seyfert galaxies NGC 2992 and 3C 120.


EGRET operations were normal this week. The percentage of possible
data that is being recovered dropped slightly this week, but is
still well above what was being recovered before the GRTS station
was in place. Preliminary analyses and delivery of preliminary
results to the SSC and Guest Investigators is now on schedule and
expected to be slightly ahead of schedule in the near future.
Interaction with guest investigators remains at a good level.

The EGRET team wishes to express their thanks to the TDRS, GRTS
station teams for a job well done--on schedule ad within budget.
The increased data return will significantly enhance the scientific
return of EGRET, and we are sure, of the Compton Observatory as a


The COMPTEL instrument is performing well and continues routine

The COMPTEL team has forwarded to the CGRO public archive at
Goddard all low-level and first high-level data products for
Viewing Periods 21 through 28 of Phase 1. The collaboration
continues to prepare for public release the remainder of such data
products for the balance of CGRO Phase 1 observations.


For the past three weeks hard x-ray flux from Cygnus X-1 has
averaged about .03-.04 photons/cm2-sec in the 45-140 keV band. 
This is the lowest flux level observed by BATSE since the beginning
of the mission.

Pulsations with a period of 103.4 s from the pulsar A0535+26 are
currently  being detected by BATSE.  The source has also been
detected with the occultation technique. The pulsar is a member of
a Be/X-ray binary system  that is near periastron passage at this
time. The pulsed flux is approximately  50 mCrab (Nebula) 20-50
keV.  Pulsed flux has been seen from the pulsar  for the last four
periastron passages.

The first delivery of FITS formatted BATSE DISCLA and CONT data was
made to the science support center last week, and is expected to be
available soon. The data delivered covers the 74 days from April
23rd to July 5th 1991, The files contain LAD count rates for all
eight detectors with either 4 energy channels and 1.024 s
resolution, or 16 energy channel an 2.048 s resolution. In addition
the files contain information on the spacecraft position and
orientation, and energy calibration data.   

BATSE folded-on-board pulsar data is currently being collected from
the Crab pulsar, PSR 1055-52, Cen X-3, and Her X-1. 

As of February 2nd, BATSE has detected 892 cosmic gamma-ray bursts
out of a total of 2692 on-board triggers in 1016 days of operation.
There have been 728 triggers due to solar flares with emission
above 60 keV.